# LanguageBind: Extending Video-Language Pretraining to N-modality by Language-based Semantic Alignment

### v1

@@ 7 Conclusion
In this work, we propose the LanguageBind, a language-based semantic alignment method for multi-modal pretraining. We employ contrastive learning to establish modality semantic alignment between the language modality and all other modalities. To improve modal integrity, we also construct the first large-scale multi-modal dataset directly aligned to language modality, VIDAL-10M, comprising 10 million aligned VL, IL, DL, and AL pairs. Experimental results on the zero-shot VL, IL, DL, and AL understanding tasks demonstrate the capability of LanguageBind and the effectiveness of VIDAL-10M.
@@ Appendix B Pretraining details
tablepretraining setting Config Video Infrared Depth Audio Vision encoder ViT-Base/32 ViT-Huge/14 ViT-Huge/14 ViT-Huge/14 Optimizer BertAdam AdamW AdamW AdamW Optimizer Momentum $\beta_{1},\beta_{2}=0.9,0.98$ Epochs 1 1 1 8 Learning rate 1e-4 1e-4 5e-4 5e-4 Coefficient learning rate 1e-3 Weight decay 0.2 Batch size 512 1024 1024 512 Warmup steps - 200 200 200 Temperature learnable Learning rate schedule cosine decay Slice frame position cut from head frames - - - Max words 32 77 77 77 Max frames 12 - - - Cls header mean-pooling - - - LoRA rank - 2 2 8 LoRA alpha - 16 16 16 LoRA dropout - 0.1 0.1 0.1
@@
For the video-text retrieval, we verify that the VIDAL-10M dataset is highly aligned. We adopted the training framework of CLIP4Clip, and the model is initialized from ViT-B/32, and the rest of the parameters are the same as the default settings, except for 1 epoch and batch size of 512.
@@
The model is initialized from OpenCLIP with a frozen language encoder. For each individual sample, we employ a random selection approach to extract either a depth image from the video sequence. Subsequently, we resize these frames to have a short edge length of 256 units, followed by a central cropping process to attain dimensions of 224×224. Additionally, we tripled the number of channels in both the depth image. The text templates employed for zero-shot classification are sourced from OpenCLIP, with a modification consisting of the substitution of ”photo” with ”depth photo” across all templates. This alteration yields an approximate performance gain of 1%.
@@
Following depth-language, it is worth noting that the text templates corresponding to infrared images retain the ”photo” designation, as no discernible performance improvement is observed from this particular modification.
@@
The data are preprocessed as in 3.1. Unlike depth and infrared, spectrograms differ much from the domain of conventional visual images. Therefore, it is not easy to overfit during training, so we increase the training epoch and the rank of LoRA. Additionally, we replace ”the/a photo of” with ”the/a sound of” across all templates for audio zero-shot classification.

### v7

@@ 6 Conclusion
In this work, we propose the LanguageBind, a language-based semantic alignment method for multi-modal pretraining. We employ contrastive learning to establish modality semantic alignment between the language modality and all other modalities. To improve modal integrity, we also construct the first large-scale multi-modal dataset directly aligned to language modality, VIDAL-10M, comprising 10 million aligned VL, IL, DL, and AL pairs. Extensive experimental results, including zero-shot X-language comprehension and indirect alignment between different modalities, demonstrate the effectiveness of LanguageBind’s multimodal alignment and complementary capabilities, as well as the effectiveness of VIDAL-10M.
@@ Appendix B Pretraining details
tableTraining setting. CLIP4Clip LanguageBind Config Video Video Infrared Depth Audio Vision encoder ViT-Base/32 ViT-Large/14 Optimizer BertAdam AdamW Optimizer Momentum $\beta_{1},\beta_{2}=0.9,0.98$ $\beta_{1},\beta_{2}=0.9,0.98$ Epochs 1 16 1 1 8 Learning rate 1e-4 1e-4 1e-4 5e-4 5e-4 Coefficient learning rate 1e-3 1 1e-3 1e-3 1e-3 Weight decay 0.2 0.2 Batch size 512 640 1024 1024 512 Warmup steps 0 2000 200 200 2000 Temperature learnable learnable Learning rate schedule cosine decay cosine decay Max words 32 77 Max frames 12 8 - - - Mask ratio - 0.3 0.5 0.5 0.3 LoRA rank - 16 2 2 16 LoRA alpha - 16 LoRA dropout - 0.0 0.1 0.1 0.1
@@
For the video-text retrieval based CLIP4Clip, we verify that the VIDAL-10M dataset is highly aligned. We adopted the training framework of CLIP4Clip, and the model is initialized from ViT-B/32, and the rest of the parameters are the same as the default settings, except for 1 epoch and batch size of 512. For the video-text retrieval based LanguageBind, we add a temporal attention before each spatial attention following Aim (Yang et al., 2023). The temporal attention is initialized from the spatial attention and LoRA is applied only to the temporal attention. We add temporal position embedding before each temporal attention. We show the details of results as shown in Table 11, Table 12 and Table 13. For zero-shot video classification, The text templates are sourced from OpenCLIP, with a modification consisting of the substitution of “photo” with “video” across all templates.
@@
The model is initialized from OpenCLIP with a frozen language encoder. For each individual sample, we employ a random selection approach to extract either a depth image from the video sequence. Subsequently, we resize these frames to have a short edge length of 256 units, followed by a central cropping process to attain dimensions of 224×224. Additionally, we tripled the number of channels in both the depth image. The text templates employed for zero-shot classification are sourced from OpenCLIP, with a modification consisting of the substitution of “photo” with “depth photo” across all templates. This alteration yields an approximate performance gain of 1%.
@@
Following depth-language, it is worth noting that the text templates corresponding to infrared images retain the “photo” designation, as no discernible performance improvement is observed from this particular modification.
@@
The data are preprocessed as in 3.1. Unlike depth and infrared, spectrograms differ much from the domain of conventional visual images. Therefore, it is not easy to overfit during training, so we increase the training epoch and the rank of LoRA. Additionally, we replace “the/a photo of” with “the/a sound of” across all templates for audio zero-shot classification.
Table 10: Zero-shot Video-Text Retrieval Performance based LanguageBind (LoRA) and 3M video-text pairs. We train it with large model. We show the details of results.
Text-to-Video Video-to-Text
Dataset R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$ R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT 42.6 65.4 75.5 2.0 37.9 63.1 73.3 3.0
MSVD 52.2 79.4 87.3 1.0 68.4 91.7 96.4 1.0
ActivityNet 35.1 63.4 76.6 3.0 32.3 62.2 74.5 3.0
DiDeMo 37.8 63.2 73.4 3.0 37.6 63.7 73.3 3.0
Table 11: Zero-shot Video-Text Retrieval Performance based LanguageBind (full tuning) and 3M video-text pairs. We train it with large model. We show the details of results.
Text-to-Video Video-to-Text
Dataset R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$ R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT 42.7 67.1 77.0 2.0 39.7 63.9 73.8 3.0
MSVD 53.5 80.5 87.5 1.0 68.1 89.5 96.0 1.0
ActivityNet 36.9 65.1 77.2 3.0 33.8 64.0 76.1 3.0
DiDeMo 38.1 65.0 73.6 2.0 38.4 63.0 72.6 3.0
Table 12: Zero-shot Video-Text Retrieval Performance based LanguageBind (full tuning) and 10M video-text pairs. We train it with large model. We show the details of results.
Text-to-Video Video-to-Text
Dataset R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$ R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT 42.8 67.5 76.0 2.0 38.3 64.0 74.1 3.0
MSVD 54.1 81.1 88.1 1.0 69.7 91.8 97.9 1.0
ActivityNet 38.4 66.6 77.9 2.0 35.7 65.8 77.8 3.0
DiDeMo 39.7 65.5 73.8 2.0 38.6 65.6 74.3 2.0
Table 13: Zero-shot Video-Text Retrieval Performance based LanguageBind (full tuning) and 10M video-text pairs. We train it with huge model. We show the details of results.
Text-to-Video Video-to-Text
Dataset R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$ R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT 44.8 70.0 78.7 2.0 40.9 66.4 75.7 2.0
MSVD 53.9 80.4 87.8 1.0 72.0 91.4 96.3 1.0
ActivityNet 41.0 68.4 80.0 2.0 39.1 69.8 81.1 2.0
DiDeMo 39.9 66.1 74.6 2.0 39.8 67.8 76.2 2.0
@@ Appendix C Downstream datasets
Audio-language.
We validate the zero-shot classification capability with the ESC-50 (Piczak, 2015) dataset, which has 2000 test audios, each uniquely labelled. For zero-shot retrieval, we use the Clotho (Font et al., 2013) dataset. Each audio has 5 corresponding captions, so we use text-to-audio retrieval to validate the model performance. We perpare test data following ImageBind.
@@ Appendix D License
Appendix E Additional Ablation Study
E.1 Impact of different text sources
In Table E.1, we conduct various experiments to explore how different text sources impact language modality. We verify the effectiveness of LanguageBind, trained with text from multiple sources, across various modalities. While some text sources yield good results, we discover that a single text source may not be universally suitable for all downstream tasks and datasets. In terms of video and depth modalities, the ChatGPT enhanced caption proves to be advantageous. For infrared images, the OFA performs best in the LLVIP dataset, while the raw caption achieves the highest accuracy in FLIR v1 and v2. That’s why our VIDAL-10M provides multi-view textual descriptions, allowing for flexibility in selecting an appropriate text source that caters to diverse task requirements.
[…]
E.2 Scaling the size of dataset
Figure 12: Scaling pretraining data size.